a computer software retailer used a markup rate of 30.find the selling price of a computer game that cost the retailer $35


1
Expert's answer
2022-02-14T08:43:10-0500

Markup = % of cost price "\times100" = "\frac{30}{100}\times 35=10.5"

Selling price = Cost price + Markup = 35 + 10.5 = $45.5